OpenStudy (anonymous):

A cylindrical water tank has a radius of 6 feet and a height of 20 feet. A larger tank with the same radius has a volume that is 2 times the volume of the smaller tank. What is the height of the larger tank?

OpenStudy (anonymous):

an easy way to remember this cylinder formula is this! you know the area of a circle right?

OpenStudy (anonymous):

no my teacher is dumb

OpenStudy (anonymous):

the area for a circle is \[\pi r ^{2}\] how many circles would it take to fill the tank 20 feet high? 20 feet worth of course! soooo for a cylinder V= \[\pi r ^{2}h\]

OpenStudy (anonymous):

first we need to find the volume of the first one. we have the radius and we have the height right!? so lets plug em in!

OpenStudy (anonymous):

\[v = \pi 6^{2} * 20\]v=pi r^2 *h

OpenStudy (anonymous):

solve that, then double that value and remember that we already have the radius so we need to solve for h